OpenStudy (anonymous):

Calculate the area of triangle ABC with altitude CD, given A (6, -2), B (1, 3), C (5, 5), and D (2, 2). (5 points) Select one: a. 12 square units b. 13 square units c. 13 square units d. 15 square units

OpenStudy (anonymous):

@timo86m @iambatman

OpenStudy (anonymous):

base * height

OpenStudy (anonymous):

what that makes no sense

OpenStudy (anonymous):

^ thats the area of a rectangle not a triangle, triangle is 1/2bh

OpenStudy (anonymous):

oh yeah, base * height * 1/2

OpenStudy (anonymous):

where is the base and the height

OpenStudy (anonymous):

Sorry guys I'm helping someone else out right now, I'll get back to you later, if you still need help.

jimthompson5910 (jim_thompson5910):

Here is what the triangle ABC looks like when you plot it on an xy coordinate grid (see attached)

jimthompson5910 (jim_thompson5910):

so you can see that AB is the base and CD is the height

OpenStudy (jdoe0001):

\(\bf \textit{Area of a Triangle}\cfrac{base\cdot height}{2}\implies \cfrac{AB\cdot CD}{2}\\ \quad \\ recall \\ \quad \\ \textit{distance between 2 points}\\ \quad \\ \qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2} \)

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&({\color{red}{ 6}}\quad ,&{\color{blue}{ -2}})\quad B&({\color{red}{ 1}}\quad ,&{\color{blue}{ 3}})\\ C&({\color{red}{ 5}}\quad ,&{\color{blue}{ 5}})\quad D&({\color{red}{ 2}}\quad ,&{\color{blue}{ 2}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)